OpenStudy (anonymous):

Input in standard form the equation of the given line. The line with m = -1 /2 and b = 1

OpenStudy (saifoo.khan):

y = mx + b Where m is the slope and b is the y-intercept, \[y = -\frac12x + 1\] \[y + \frac12x = 1\]

OpenStudy (anonymous):

Standard form is Ax+By=C

OpenStudy (saifoo.khan):

That's the same as i wrote.

OpenStudy (anonymous):

the answer shouldn't have fractions in it though

OpenStudy (saifoo.khan):

1/2 = 0.5

OpenStudy (anonymous):

there are also no decimals

OpenStudy (saifoo.khan):

Multiply the whole equation by 2.